The size of the labor force is 400 million. Each month 36% of unemployed workers find jobs and 4% of employed workers lose their jobs.

How many unemployed people find jobs each month (in millions)
Answer #1

36% of unemployed workers find jobs every month.

So, f = 36% = 0.36

4% of employed workers lose their jobs.

So, s = 4% = 0.04

Steady state unemployment rate = s / (s+f)

Steady state unemployment rate = 0.04 / (0.04 + 0.36)

Steady state unemployment rate = 0.1

Steady-state unemployment rate = U / L

Where U is unemployed workers.

L is labor force.

0.1 = (U / 400 million)

U = 0.1 (400 million)

U = 40 million.

Unemployed workers = 40 million.

36% of unemployed workers find jobs each month.

Number of unemployed people finds jobs each month = 36% of 40 million

Number of unemployed people finds jobs each month = 0.36 (40 million)

The number of unemployed people finds jobs each month = 14.4 million.